OpenStudy (anonymous):

Anton joined a golf club two years ago. He pays an annual membership fee of $895 and a greens fee of $30 each time he plays a game of golf. The function below can be used to calculate the total yearly golfing fee, f (g), in dollars. f (g) = 895 + 30g, where g represents the number of times he played golf during the year. Last year he paid $2,065 as a total golfing fee. For how many games did he pay a greens fee ?

OpenStudy (anonymous):

895 + 30g = 2,065 => g = (2.065 - 895)/ 30 = 39 games

OpenStudy (anonymous):

thank you soo much you really helped